UT Feb 27 2021




Night Partition

  • Night split with LBTI - night center 00:32 - handover 00:27 to LBTI
  • Moon is up effectively all night (98.4%)
  • 1st half D/E-time
    • DONETwilight (19:15hr)-21:15 - HD 33632 - LBTI w/LMIRCAM - M-Band Imaging
    • DONELBCs - TMS testing
  • 2nd half LBTI
NO LUCI-1 SOUL AO

Night Plan for D/E time

  • Engineering Tasks
    • LBCs TMS Testing
      • Targets TBD
  • Science
    • HD33632 brown dwarf companion observation - LBTI/LMIRCAM - (Stone) - 2hr
      • 05 13 17 +37 20 14
      • Directors time from December, lost to weather and M3 issue
      • Target from twilight + 2hrs
      • Weather: clear, <1.2-ish arcsec seeing

Overview

20210226 23:54 Powering up LBC's and prepping directories. Powering up LBTI WFS, and getting all other toys ready for tonight.

20210227 0:42 David has complted his test preset and telescope checkout. Taking some LBC darks. Taking the MODS calibs from yesterday too.
  • simSnap & 1.2" flats for MODs
  • biases for LBC. Issues wit
  • RB_Science on robs dies with new image and wont dispay (IT8329)
20210227 6:20 David is slewing to 30 and opening.

See LBTI log for additional details: https://docs.google.com/spreadsheets/d/1NLjq6WSJSXHLfO8VKsPOhHgqU2T5KFzDFRHNqQglnxs/edit?usp=sharing

20210227 1:35 Setup on Capella. There was an active flao preset left from daywork. I thought by cancelling it would clear it but it was left there during authorization and it is super stiicky. WFS spec keeps reveryting to FLAOWFS and failing our preset. We are reauthorizing. That fixed it. TN data taken 014445

20210227 1:49 Slew to HD33632.OCAM not live, change of binning resolved. Sweetspots marked. TN#015926 There was water vapor fingers in spectrum so abandoning ALES and focusing on imaging with LBTI time tonight.

20210227 2:05 Taking M band data!

20210227 7:41. The telescope is complang of a failed offset on our side...but we didn't offset and the loop is still closed and we are still taking data. So not sure why. But there is a big "Failed" in IIF for our offset? TN# 024204

Steve is manually sending an offset to verify it works. (0,0) ok, and (-12,0) worked as expected. Continuing with script. Maybe we took too long to close. Seeing 0.83"

20210227 3:31 Negative average pupil intensity after offset. Loop opened. Rebinning resolved.

20210227 3:55 Reoptimizing a few times...PSF less stable than earlier. Manually optimized and looks better

20210227 4:10 Taking unsats then done. David has lost the DIMM and is working to recover.

20210227 4:14 Reconfig from LBTI/PEPSI to LBC's

20210227 4:22 TMS lasers on and stabilized

20210227 4:37 David had issues with the reconfigure on the right. The right arms didn't move. Readjusting

20210227 4:53 Reconfig complete. Olga has connected LBC's. Moving to first field ~70deg elevation for focus and collimation. running do hybrid. The sky is too bright. Could not find pupils. Collimating on first test field at 57 degrees (a little further from the moon)

From Olga's Log: https://docs.google.com/document/d/1I-FY7vkzHfqwzR_lyvdhJ3Tc0PfOogI_kG_8z4YESHw/edit?usp=sharing

Test 1: TMS active corrections on in-focus field

05:01 Going to the field for the first test that is a little lower elevation. The initial z4/z11 from our first field might have been left in and I didn't back it out. David is clearing active optics.

05:03 dohybrid. Did not converge, used "S" to stop. Z22 was building up.

05:16 Setting the reference.

/lbt/lbto/supportscripts/MetrologySupport/lbc_set_tms_offset.py -s 1

Had to clear an existing file

Tried again and it was successful. Took 9 seconds.

05:19 Started a script to take a series of 99 Blue (V) and Red (r-SLOAN) images

05:20 Started TMS active:

A problem: **=> initiating a TMS reading... measurement failed (1.096s)

This is a known issue having to do with MATLAB and the Etalon SW. TCP error - Andrew says once

this happens, one has to restart the Etalon SW.

05:28 Andrew restarted Etalon SW & TCP.

05:29 Restarted the LBC script and tms loop

LBCR 052936 has a jump (it replicated the pointing offsets).

05:39 John says that since we started -6 mic of Z have been applied.

Glass = -0.36, Ambient = -1.128. When we started it was about Glass/Ambient 0/-0.8.

06:03 Stopped TMS and LBC script

Test 2: Series of pupil images with TMS running in active mode

06:03 New test:

dofpia, RedLimit =1, BlueLimit =1

Set Reference after first set of corrections have been sent, so as not to collide:

lbc__set_tms_offset.py -s 1 (did not get to this step because of runaway)

06:08: dofpia was running away in Z11/Z22. I control C'd and David is clearing active optics.

06:10 dofpia to converge (or try...) Glass/Ambient = -0.7/-1.2 C

Shift "S" to end.

06:14 Set reference: lbc_set_tms_offset.py -s 1

David is adding -875 nm Z4 and +500 nm Z11 to DX

Set reference --- oops, Z11 was cleared in adding Z4.

06:19 Now we are really adding -875 Z4 and +500 nm Z11 to DX and

I am setting the TMS reference.

06:21 TMS loop started (tms_loop_rightonly.pl) and RB_Vlowextra99 script started.

Error: Added the pointing offset on again.

06:47 I noticed an error in the restarted the RB_Vlowextra99 script as I had changed the 2nd NExpo from 1 to 99 in the wrong xml tag (there are 8 per channel and I must have changed two in the Blue and none in the Red). So we had one LBCR pupil image: lbcr.20210227.062125.fits and the next one will be lbcr.20210227.064723.fits

Test 3: Add Z11 to see if TMS SW will correct it.

07:10 John is changing Z11 by +500nm. Only changed by 2 microns.

07:13 John is sending Z22 by +100nm. Measurement failed (Etalon TCP error).

7:15 ok and on that note..., we are done. Going up to zenith and reconfiguring for LBTI/PEPSI.

7:25 TMS lasers off

7:40 We are off to Beta Leo, first LBTI target

7:50 RIP on center pupils. We were pretty aberrated when ran it and not all that well centered on cross hairs. Recover fail at el 68, set flat with 639 modes. Fixed up a little more and we were off to the races.

-- BarryRothberg - 25 Feb 2021
Topic revision: r6 - 28 Feb 2021, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback